# Stable Offline Value Function Learning with Bisimulation-based Representations

### v3

@@ 4.3 Offline Policy Evaluation
In this set of experiments, we evaluate the algorithms for ope on $13$ datasets: $4$ dm control and $9$ d4rl datasets. We also evaluate bcrl-na, which is bcrl without the exploration maximization regularizer. To stabilize training for all algorithms, we use wide neural networks with layernorm (Gallici et al., 2024; Ota et al., 2021). Note that while wide networks and layernorm stabilize training, they may not lead to stable lspe under the linear evaluation protocol. We report the final (normalized) squared value errors when the learned representations are used with lspe for ope. The results are presented in Table 1 (see Appendix C.3.2 for full learning curves).
@@
In general, we find that krope representations lead to low and stable msve. On the other hand, we find that the other auxiliary tasks inconsistently produce stable ope estimates across all environments. Compared to the other bisimulation-based algorithms (dbc and rope), krope consistently achieves lower error. In all cases, dbc seems to be unreliable for ope. rope, which was the previous state-of-the-art bisimulation-based algorithm for ope, also achieves low error. While rope is competitive with krope, krope generally outperforms rope and has one less hyperparameter. The performance of the abs-dr3 and beer regularizer suggests that explicitly trying to increase the rank of the features of the penultimate layer may hurt stability, and even if the ope error is stable, it can hurt accuracy. We also make a similar observation for bcrl. However, in this case, we attribute poor performance to difficulty in optimizing the bcrl objective. In fact, in Figure 3(a), we will see that bcrl is sensitive to hyperparameter tuning. We also observe results consistent with a known result that bcrl-exp-na achieves lower error than bcrl-na indicating the known result that exploration maximization of the covariance matrix helps produce stable representations (Chang et al., 2022). While fqe achieves low error on WalkerStand and CheetahRun, it is very unstable on the other datasets, which motivates the need to shape the representations for stable and accurate ope. Note that in practice, krope may still diverge, as it did it in $3$ cases, because the learning objective in Equation (4) is still a semi-gradient learning algorithm (see discussion in Section 5 and Appendix C.3.1).
@@
Takeaway #$\#2$: Practical Stable and Accurate Offline Policy Evaluation ope practitioners can use krope for stable and accurate evaluation of offline rl agents.
@@ 5 Limitations and Future Work
In this section, we discuss limitations and future work. A shortcoming of our work is that krope’s learning algorithm is susceptible to instability since it is a semi-gradient method (Sutton & Barto, 2018). Moreover, its fixed-point optimization means it does not solve any objective function (Feng et al., 2019). In our work, we employed commonly-used techniques such as layernorm and wide neural networks to mitigate instability (Ota et al., 2021; Gallici et al., 2024). While these techniques potentially side-step the issue, the consequences of a semi-gradient method may still exist. In Appendix C.3.1, we present an empirical analysis to gauge when krope’s learning algorithm may be unstable. We find that while individual off-policy transitions can determine the instability of fixed-point and semi-gradient algorithms such as fqe, pairs of off-policy transitions can determine krope’s instability. Since we are unlikely to have control over the distribution over pairs of transitions in practice, we need to resort to fundamental changes to the algorithm. One potential modification to overcome this limitation is based on that by Feng et al. (2019). Their key insight is to leverage the Legendre-Fenchel transformation and replace the fixed-point loss function of semi-gradient methods with an equivalent expression that avoids semi-gradient learning (Rockafellar & Wets, 1998). However, a drawback with this approach is that the new learning objective is a minimax procedure, which can be challenging to optimize in practice. In future work, we will explore the viability of this approach to design a provably convergent version of krope.
@@ 6 Conclusion
In this work, we tackled the problem of stabilizing offline value function learning in reinforcement learning. We introduced a bisimulation-based representation learning algorithm, kernel representations for ope (krope), that shapes the state-action representations to stabilize this procedure. Theoretically, we showed that krope representations are stable from two perspectives: 1) non-expansiveness, i.e., they lead to value function learning updates that enable convergence to a fixed-point and 2) Bellman completeness, i.e., they satisfy a condition for data-efficient policy evaluation. Empirically, we showed that krope leads to more stable and accurate offline value function learning than baselines. Our work showed that bisimulation-based representation learning effectively stabilizes long-term performance evaluations of offline reinforcement learning agents.
@@ A.1 Bisimulation Metrics
In this section, we present background information on bisimulations and its associated metrics. Our proposed representation learning algorithm is a bisimulation-based algorithm. Bisimulation abstractions are those under which two states with identical reward functions and that lead to identical groups of next states under any action are classified as similar (Ferns et al., 2004, 2011; Ferns & Precup, 2014). Bisimulations are the strictest forms of abstractions. In practice, the exact bisimulation criterion is difficult to satisfy computationally and statistically. A more relaxed version of this notion is the $\pi$-bisimulation metrics. These metrics capture the similarity between two states such that two states are considered similar if they have identical expected reward functions and expected transitions to identical groups of next states under a fixed policy $\pi$ (Castro, 2019).
@@
(Castro, 2019) An abstraction $\phi:\mathcal{X}\to\mathcal{X}^{\phi}$ over the state-action space $\mathcal{X}$ is a $\pi$-bisimulation for a fixed policy $\pi$ if for any two state-actions $x,y\in\mathcal{X}$ and abstract state-action $x^{\phi}\in\mathcal{X}^{\phi}$, $\phi$ is such that $\phi(x)=\phi(y)$ implies that $r(x)=r(y)$ and $\sum_{x^{\prime}\in x^{\phi}}P^{\pi}(x^{\prime}|x)=\sum_{x^{\prime}\in x^{\phi}}P^{\pi}(x^{\prime}|y)$.
@@
(Castro, 2019) Let $\mathcal{M}(\mathcal{X})$ be the space of bounded pseudometrics on the state-action space $\mathcal{X}$ and $\pi$ be a fixed policy. Then define $\mathcal{B}:\mathcal{M}(\mathcal{X})\to\mathcal{M}(\mathcal{X})$ such that for each $d\in\mathcal{M}(\mathcal{S})$:
@@
Using the above metrics, prior works have introduced several representation learning algorithms to learn representations such that the distance between representations in latent space model the above distance metrics (Castro et al., 2022, 2023; Zhang et al., 2021; Kemertas & Aumentado-Armstrong, 2021; Pavse & Hanna, 2023a).
@@ Appendix B Theoretical Results
Lemma 3 tells us that the krope state-actions that are close in latent space also have similar action-values upto a constant $C:=\frac{1}{2}\sum^{m+1}_{n=0}\gamma^{n}(\Delta_{n}^{{\pi_{e}}}(x)+\Delta_{n}^{{\pi_{e}}}(y))$. Intuitively, $\Delta_{n}^{{\pi_{e}}}(x)$ is the expected absolute reward difference between two trajectories at the $n$th step after ${\pi_{e}}$ is rolled out from $x$. If the transition dynamics and ${\pi_{e}}$ are deterministic, we have $C=0$ (Castro, 2019; Zhang et al., 2021). Note that while the deterministic transition dynamics assumption is eliminated, the bound suggests that krope may hurt accuracy of $\hat{q}^{{\pi_{e}}}$ since when $d_{\text{{krope}}}(x,y)=0$, we get $|q^{\pi_{e}}(x)-q^{\pi_{e}}(y)|\leq C$. This indicates that two state-actions that may have different action-values are considered the same under krope. This implies that while $x$ and $y$ should have different representations, they actually may have the same representation.
@@ B.2.1 Supporting Theoretical Results
We first define ${\pi_{e}}$-bisimulation (Castro, 2019). Note that (Castro, 2019) considered only state abstractions, while we consider state-action abstractions. $\phi$ is considered a ${\pi_{e}}$-bisimulation abstraction if it induces a mapping between $\mathcal{X}$ and $\mathcal{X}^{\phi}$ such that for any $x,y\in\mathcal{X}$ such that $x,y\in\phi(x)$, we have:
@@ C.1 Empirical Setup
In all the continuous state-action experiments, we use a neural network with $1$ layer and $1024$ neurons using relu activation function and layernorm to represent the encoder $\phi:\mathcal{X}\to\mathbb{R}^{d}$ (Gallici et al., 2024). We use mini-batch gradient descent to train the network with mini-batch sizes of $2048$ and for $500$ epochs, where a single epoch is a pass over the full dataset. We use the Adam optimizer with learning rate $\{1e^{-5},2e^{-5},5e^{-5}\}$ and weight decay $1e^{-2}$. The target network is updated with a hard update after every epoch. The output dimension $d$ is $\{|\mathcal{X}|/4,|\mathcal{X}|/2,3|\mathcal{X}|/4\}$, where $|\mathcal{X}|$ is the dimension of the original state-action space of the environment. All our results involve analyzing this learned $\phi$. Since fqe outputs a scalar, we add a linear layer on top of the $d$-dimensional vector to output a scalar. The entire network is then trained end-to-end. The discount factor is $\gamma=0.99$. The auxiliary task weight with fqe for all representation learning algorithms is $\alpha=0.1$. When using lspe for ope, we invert the covariance matrix by computing the pseudoinverse.
In the tabular environments, we use a similar setup as above. The only changes are that we use a linear network with a bias component but no activation function and fix the learning rate to be $1e^{-3}$. For experiment in Appendix C.3.1, $\alpha=0.8$. We refer the reader pseudo-code in Appendix A.

### v4

@@ 4.3 Offline Policy Evaluation
In this set of experiments, we evaluate the algorithms for ope on $13$ datasets: $4$ dm control and $9$ d4rl datasets. We also evaluate bcrl-na, which is bcrl without the exploration maximization regularizer. To stabilize training for all algorithms, we use wide neural networks with layernorm (Gallici et al., 2025; Ota et al., 2024). Note that while wide networks and layernorm stabilize training, they may not lead to stable lspe under the linear evaluation protocol. We report the final (normalized) squared value errors when the learned representations are used with lspe for ope. The results are presented in Table 1 (see Appendix C.3.1 for full learning curves).
@@
In general, we find that krope representations lead to low and stable msve on $10/13$ datasets. On the other hand, we find that the other auxiliary tasks inconsistently produce stable ope estimates across all environments. Compared to the other bisimulation-based algorithms (dbc and rope), krope consistently achieves lower error. In all cases, dbc seems to be unreliable for ope. rope, which was the previous state-of-the-art bisimulation-based algorithm for ope, also achieves low error. While rope is competitive with krope, krope generally outperforms rope and has one less hyperparameter. The performance of the abs-dr3 and beer regularizer suggests that explicitly trying to increase the rank of the features of the penultimate layer may hurt stability, and even if the ope error is stable, it can hurt accuracy. We also make a similar observation for bcrl. However, in this case, we attribute poor performance to difficulty in optimizing the bcrl objective. In fact, in Figure 3(a), we will see that bcrl is sensitive to hyperparameter tuning. We also observe results consistent with a known result that bcrl-exp-na achieves lower error than bcrl-na indicating the known result that exploration maximization of the covariance matrix helps produce stable representations (Chang et al., 2022). While fqe achieves low error on WalkerStand and CheetahRun, it is very unstable on the other datasets, which motivates the need to shape the representations for stable and accurate ope. Note that in practice, krope may still diverge, as it did it in $3$ cases, because the learning objective in Equation (4) is still a semi-gradient learning algorithm (see discussion in Section 5 and Section 4.4).
@@
Takeaway #$\#2$: Practical Offline Policy Evaluation with krope krope can increase the stability and accuracy of evaluation of offline rl agents.
4.4 Stability of the krope Learning Procedure
In this section, we compare the susceptibility of krope’s and fqe’s semi-gradient learning to divergence. We refer the reader to Appendix C.1 for more details.
We conduct our experiments on the Markov reward process (mrp) in Figure 4(a) (see Feng et al. (2019)). The mrp consists of $4$ non-terminal states, $1$ terminal state (the box), and only $1$ action. The value function estimate is linear in the weights $w=[w_{1},w_{2},w_{3}]$, so, starting clockwise from the left-most circle, the native features of the states are $[1,0,0]$, $[0,1,0]$ , $[0,0,2]$, and $[0,0,1]$. In this setup, we say a transition is bad if the bootstrapping target is a moving target for the current state. For example, the transition from $w_{3}$ to $2w_{3}$ is a bad transition since updates made to $w_{3}$ may move $2w_{3}$ further away. When this transition is sampled at a frequency that is different from the on-policy distribution, algorithms such as td, lspe, and fqe tend to diverge (Asadi et al., 2024). Similarly, for krope, which uses its weights to compute the latent representation of states, we would expect that pairs of transitions that lead to moving dot product targets are bad transition pairs.
(a) Divergence Counterexample
(b) krope mse Loss
(c) fqe+krope mse Loss
Figure 4: Figure 4(a): mrp counterexample designed to illustrate divergence; $r$ denotes the rewards and $p$ denotes the probability of transition (Feng et al., 2019). Figures 4(b) and 4(c): krope training loss and fqe+krope training loss vs. epochs respectively when different datasets are paired with $\mathcal{D}_{1}$; results are averaged over $20$ trials, shaded region is the $95\%$ confidence interval, and lower is better.
To understand the stability of the learning procedures, we design the following experiment when using krope as an auxiliary loss function to fqe. We first create an on-policy dataset $\mathcal{D}^{\text{on}}$. We then define off-policy datasets of the form $\mathcal{D}^{s}$, which consists of transitions starting from the specified state $s$, where $s=\{w_{1},w_{2},w_{3},2w_{3}\}$. Using these datasets, we then construct $\mathcal{D}_{1}=\mathcal{D}^{\text{on}}\cup\mathcal{D}^{w_{3}}$ and four $\mathcal{D}_{2}$ variations: 1) $\mathcal{D}_{2}^{w_{1}}=\mathcal{D}^{\text{on}}\cup\mathcal{D}^{w_{1}}$, 2) $\mathcal{D}_{2}^{w_{2}}=\mathcal{D}^{\text{on}}\cup\mathcal{D}^{w_{2}}$, 3) $\mathcal{D}_{2}^{w_{3}}=\mathcal{D}^{\text{on}}\cup\mathcal{D}^{w_{3}}$, and 4) $\mathcal{D}_{2}^{2w_{3}}=\mathcal{D}^{\text{on}}\cup\mathcal{D}^{2w_{3}}$. Therefore, $\mathcal{D}_{1},\mathcal{D}_{2}^{w_{3}},\mathcal{D}_{2}^{2w_{3}}$ consist of mostly bad transitions while $\mathcal{D}_{2}^{w_{1}},\mathcal{D}_{2}^{w_{2}}$ consist of good transitions. Since krope samples pairs of transitions, we investigate the consequences of pairing $\mathcal{D}_{1}$ with a $\mathcal{D}_{2}$ variation. We pair each of the four $\mathcal{D}_{2}$ datasets with $\mathcal{D}_{1}$ and train fqe+krope, where fqe is fed transition samples from $\mathcal{D}_{1}$ and krope is fed samples from $\mathcal{D}_{1}$ and the specific $\mathcal{D}_{2}$ variation.
In Figure 4(b), we show the training loss of only krope to analyze when krope may diverge. We find that even though $\mathcal{D}_{1}$ consists of mostly bad transitions, if krope samples good transitions from $\mathcal{D}_{2}^{w_{1}}$ or $\mathcal{D}_{2}^{w_{2}}$, its divergence is mitigated. However, if krope samples bad transitions from $\mathcal{D}_{2}^{w_{3}}$ or $\mathcal{D}_{2}^{2w_{3}}$, it diverges since the pairing of samples from $(\mathcal{D}_{1},\mathcal{D}_{2}^{w_{3}})$ and $(\mathcal{D}_{1},\mathcal{D}_{2}^{2w_{3}})$ leads to krope chasing a moving bootstrapped dot product target. Therefore, sampling pairs of bad transitions can make krope more likely to diverge, while sampling a single bad transition (with a good transition) can make it less likely to diverge.
In Figure 4(c), we plot the combined training loss of fqe+krope. As expected, when fqe uses only bad transitions from $\mathcal{D}_{1}$, it diverges (No Pairing). In fact, in all cases, the divergence is due to fqe even when the corresponding krope variation is not diverging (see $\mathcal{D}_{2}^{w_{1}}$ or $\mathcal{D}_{2}^{w_{2}}$ in Figure 4(b)). For the $\mathcal{D}_{2}^{w_{3}}$ or $\mathcal{D}_{2}^{2w_{3}}$ variations, both fqe and krope diverge. Therefore, sampling single bad transitions makes fqe more likely to diverge (Asadi et al., 2024).
While we hand-designed these datasets, in general, we would expect that the probability of sampling a pair of bad transitions is less than that of sampling a single bad transition [^1] We expect this statement to roughly hold true unless the data consists of mostly bad transitions; see Appendix C.1. These experiments show that while value function learning with fqe may diverge, representation learning with krope may not.
Takeaway #$\#3$: krope Divergence While fqe’s divergence is due to sampling bad transitions, krope’s divergence is due to sampling bad pairs of transitions. Intuitively, since the probability of sampling a bad transition pair is less than that of sampling a single bad transition, krope training may be easier than fqe training.
@@ 5 Limitations and Future Work
In this section, we discuss limitations and future work. A shortcoming of krope’s semi-gradient algorithm is its susceptibility to divergence (Sutton & Barto, 2018). While we employed commonly-used techniques such as layernorm and wide neural networks to mitigate divergence (Ota et al., 2024; Gallici et al., 2025), the consequences of a semi-gradient method may still exist. One potential remedy is to avoid semi-gradient learning. Feng et al. (2019) suggest to replace the fixed-point loss function of semi-gradient methods with an equivalent expression that leverages the Legendre-Fenchel transformation (Rockafellar & Wets, 1998). However, a drawback with this approach is that the new learning objective is a minimax procedure, which can be challenging to optimize in practice. In future work, we will explore the viability of this approach to design a provably convergent version of krope.
@@ 6 Conclusion
In this work, we tackled the problem of stabilizing offline value function learning in reinforcement learning. We introduced a bisimulation-based representation learning algorithm, kernel representations for ope (krope), that shapes the state-action representations to stabilize this procedure. Theoretically, we showed that krope representations are stable from two perspectives: 1) non-expansiveness, i.e., they lead to value function learning updates that enable convergence to a fixed-point and 2) Bellman completeness, i.e., they satisfy a condition for data-efficient policy evaluation. Empirically, we showed that krope leads to more stable and accurate offline value function learning than baselines. We also demonstrated when representation learning with krope may be easier than value function learning with fqe. Our work showed that bisimulation-based representation learning can improve the stability and accuracy of long-term performance evaluations of offline reinforcement learning agents.
@@
Acknowledgments
The authors thank Abhinav Narayan Harish, Adam Labiosa, Andrew Wang, Anshuman Senapati, David Abel, Kyle Domico, Ishan Durugkar, Pablo Samuel Castro, Prabhat Nagarajan, Tengyang Xie, and the anonymous reviewers at iclr and icml for feedback on our work, and thank Fanghui Liu for inspiring discussions on kernel methods. Y. Chen is partially supported by NSF CCF-2233152. Q. Xie is supported in part by National Science Foundation (NSF) grants CNS-1955997, EPCN-2339794 and EPCN-2432546. J. Hanna acknowledges support from NSF (IIS-2410981), American Family Insurance through a research partnership with the University of Wisconsin—Madison’s Data Science Institute, and Sandia National Labs through a University Partnership Award.
@@ A.1 Bisimulation Metrics
In this section, we present background information on bisimulations and its associated metrics. Our proposed representation learning algorithm is a bisimulation-based algorithm. Bisimulation abstractions are those under which two states with identical reward functions and that lead to identical groups of next states under any action are classified as similar (Ferns et al., 2004, 2011; Ferns & Precup, 2014). Bisimulations are the strictest forms of abstractions. In practice, the exact bisimulation criterion is difficult to satisfy computationally and statistically. A more relaxed version of this notion is the $\pi$-bisimulation metrics. These metrics capture the similarity between two states such that two states are considered similar if they have identical expected reward functions and expected transitions to identical groups of next states under a fixed policy $\pi$ (Castro, 2020).
@@
(Castro, 2020) An abstraction $\phi:\mathcal{X}\to\mathcal{X}^{\phi}$ over the state-action space $\mathcal{X}$ is a $\pi$-bisimulation for a fixed policy $\pi$ if for any two state-actions $x,y\in\mathcal{X}$ and abstract state-action $x^{\phi}\in\mathcal{X}^{\phi}$, $\phi$ is such that $\phi(x)=\phi(y)$ implies that $r(x)=r(y)$ and $\sum_{x^{\prime}\in x^{\phi}}P^{\pi}(x^{\prime}|x)=\sum_{x^{\prime}\in x^{\phi}}P^{\pi}(x^{\prime}|y)$.
@@
(Castro, 2020) Let $\mathcal{M}(\mathcal{X})$ be the space of bounded pseudometrics on the state-action space $\mathcal{X}$ and $\pi$ be a fixed policy. Then define $\mathcal{B}:\mathcal{M}(\mathcal{X})\to\mathcal{M}(\mathcal{X})$ such that for each $d\in\mathcal{M}(\mathcal{S})$:
@@
Using the above metrics, prior works have introduced several representation learning algorithms to learn representations such that the distance between representations in latent space model the above distance metrics (Castro et al., 2021, 2023; Zhang et al., 2021; Kemertas & Aumentado-Armstrong, 2021; Pavse & Hanna, 2023a).
@@ Appendix B Theoretical Results
Lemma 3 tells us that the krope state-actions that are close in latent space also have similar action-values upto a constant $C:=\frac{1}{2}\sum^{m+1}_{n=0}\gamma^{n}(\Delta_{n}^{{\pi_{e}}}(x)+\Delta_{n}^{{\pi_{e}}}(y))$. Intuitively, $\Delta_{n}^{{\pi_{e}}}(x)$ is the expected absolute reward difference between two trajectories at the $n$th step after ${\pi_{e}}$ is rolled out from $x$. If the transition dynamics and ${\pi_{e}}$ are deterministic, we have $C=0$ (Castro, 2020; Zhang et al., 2021). Note that while the deterministic transition dynamics assumption is eliminated, the bound suggests that krope may hurt accuracy of $\hat{q}^{{\pi_{e}}}$ since when $d_{\text{{krope}}}(x,y)=0$, we get $|q^{\pi_{e}}(x)-q^{\pi_{e}}(y)|\leq C$. This indicates that two state-actions that may have different action-values are considered the same under krope. This implies that while $x$ and $y$ should have different representations, they actually may have the same representation.
@@ B.2.1 Supporting Theoretical Results
We first define ${\pi_{e}}$-bisimulation (Castro, 2020). Note that (Castro, 2020) considered only state abstractions, while we consider state-action abstractions. $\phi$ is considered a ${\pi_{e}}$-bisimulation abstraction if it induces a mapping between $\mathcal{X}$ and $\mathcal{X}^{\phi}$ such that for any $x,y\in\mathcal{X}$ such that $x,y\in\phi(x)$, we have:
@@ B.3 Main krope Theoretical Results
Remark on the injective reward assumption. The injective reward assumption simply means that each abstract state-action group will have a distinct associated reward from every other abstract state-action group. This assumption comes as a trade-off. Chen & Jiang (2019) proved that bisimulation abstractions are Bellman complete. Instead of assuming injective rewards, they assumed that two states were grouped together if each state’s transition dynamics led to next states that are also grouped together (one of the conditions for exact bisimulations). This condition is also considered strict and inefficient to compute (Castro, 2020).
In our work, we relax this exact transition dynamics equality by considering independent couplings between next state distributions, thereby making the krope algorithm efficient to compute. However, the drawback of this relaxation is that preserving distinctness between abstract state-actions may be lost (see remarks on $k^{\pi_{e}}$ in Section 3.1). To ensure the distinctness between state-action abstractions, we assumed that the reward function is injective. This then allowed us to show Bellman completeness in a similar way to that shown in Chen & Jiang (2019).
@@ C.1 Empirical Setup
In all the continuous state-action experiments, we use a neural network with $1$ layer and $1024$ neurons using relu activation function and layernorm to represent the encoder $\phi:\mathcal{X}\to\mathbb{R}^{d}$ (Gallici et al., 2025). We use mini-batch gradient descent to train the network with mini-batch sizes of $2048$ and for $500$ epochs, where a single epoch is a pass over the full dataset. We use the Adam optimizer with learning rate $\{1e^{-5},2e^{-5},5e^{-5}\}$ and weight decay $1e^{-2}$. The target network is updated with a hard update after every epoch. The output dimension $d$ is $\{|\mathcal{X}|/4,|\mathcal{X}|/2,3|\mathcal{X}|/4\}$, where $|\mathcal{X}|$ is the dimension of the original state-action space of the environment. All our results involve analyzing this learned $\phi$. Since fqe outputs a scalar, we add a linear layer on top of the $d$-dimensional vector to output a scalar. The entire network is then trained end-to-end. The discount factor is $\gamma=0.99$. The auxiliary task weight with fqe for all representation learning algorithms is $\alpha=0.1$. When using lspe for ope, we invert the covariance matrix by computing the pseudoinverse.
In the tabular environments, we use a similar setup as above. The only changes are that we use a linear network with a bias component but no activation function and fix the learning rate to be $1e^{-3}$. For the experiment in Section 4.4, $\alpha=0.8$. We refer the reader pseudo-code in Appendix A.
@@
Toy Divergence Example
We set $\gamma=1$. The sampled state-actions are fed into a linear network encoder with a bias component and no activation function which outputs a $d=3$ representation. This representation is shaped by krope and this representation is then fed into a linear function to output the scalar value for fqe. $\mathcal{D}^{\text{on}}$ is a dataset of $2000$ on-policy transitions. $\mathcal{D}^{s}$ is an off-policy dataset with $5000$ transitions starting from the specified state $s$, where $s=\{w_{1},w_{2},w_{3},2w_{3}\}$, where the next state is sampled according to the transition dynamics of the mrp.
Regarding the statement of, “we would generally expect that the probability of sampling a bad transition pair is less than that of sampling a single bad transition”, we make the following simplistic and rough calculation. For a given a dataset, assume that the probability of choosing a bad transition is $p$ and a good transition is $1-p$. Then sampling a pair of bad transitions versus sampling a single bad transition involves comparing the probabilities of $p^{2}$ and $2p(1-p)$, respectively. In general, we would expect $2p(1-p)\geq p^{2}$, unless $p\geq 0.66$, i.e., the dataset starts to overflow with bad transitions. Roughly, we would expect fqe to diverge when $p\geq 0.5$, but we would expect krope to diverge when $p\geq 0.66$, which suggests that krope is less prone to divergence than fqe.